find the curvature and torsion of the curve z=u,y=1+u/u,z=1-u^2/u
1
Expert's answer
2022-02-01T15:16:25-0500

a)


"r(u)=\langle u, \dfrac{1+u}{u}, \dfrac{1-u^2}{u}\rangle"

"r'(u)=\langle1, -\dfrac{1}{u^2},-\dfrac{1}{u^2}-1\rangle"

"r''(u)=\langle0, \dfrac{2}{u^3}, \dfrac{2}{u^3}\rangle"

"|r'(u)|=\sqrt{(1)^2+(-\dfrac{1}{u^2})^2+(-\dfrac{1}{u^2}-1)^2}"

"=\dfrac{\sqrt{2u^4+2u^2+2}}{u^2}"

"r'(u)\times r''(u)=\begin{vmatrix}
 i & j & k \\
\\
 1 & -\dfrac{1}{u^2} & -\dfrac{1}{u^2}-1 \\ \\
 0 & \dfrac{2}{u^3}&\dfrac{2}{u^3}
\end{vmatrix}"

"=i\begin{vmatrix}
 -\dfrac{1}{u^2} & -\dfrac{1}{u^2}-1 \\ \\
 \dfrac{2}{u^3} & \dfrac{2}{u^3}
\end{vmatrix}-j\begin{vmatrix}
 1 & -\dfrac{1}{u^2}-1 \\ \\
 0& \dfrac{2}{u^3}
\end{vmatrix}"

"+k\begin{vmatrix}
 1 & -\dfrac{1}{u^2} \\ \\
 0 & \dfrac{2}{u^3}
\end{vmatrix}=-\dfrac{2}{u^3}i-\dfrac{2}{u^3}j+\dfrac{2}{u^3}k"

"|r'(t)\times r''(t)|=\sqrt{(-\dfrac{2}{u^3})^2+(-\dfrac{2}{u^3})^2+(\dfrac{2}{u^3})^2}"

"=\dfrac{2\sqrt{3}}{u^2|u|}"

Find curvature


"\kappa(t)=\dfrac{|r'(t)\times r''(t)|}{(|r'(t)|)^{3}}"

"=\dfrac{\dfrac{2\sqrt{3}}{u^2|u|}}{(\dfrac{\sqrt{2u^4+2u^2+2}}{u^2})^{3}}"

"=\dfrac{\sqrt{6}u^2|u|}{2(u^4+u^2+1)^{3/2}}"

"\kappa(u)=\dfrac{\sqrt{6}u^2|u|}{2(u^4+u^2+1)^{3/2}}"

b)


"r'''(u)=\langle0, -\dfrac{6}{u^4}, -\dfrac{6}{u^4}\rangle"

"(r'(u)\times r''(u))\cdot r'''(u)=0-\dfrac{2}{u^3}(-\dfrac{6}{u^4})+\dfrac{2}{u^3}(-\dfrac{6}{u^4})"

"=0"


"\tau(u)=\dfrac{(r'(u)\times r''(u))\cdot r'''(u)}{(|r'(t)\times r''(t)|)^2}=0"